### v3

# Milnor number of an invariant singularity: generalization of Chulkov’s inequality

Abstract.
S.P. Chulkov has proven that the Milnor number of a function germ $f$ with zero 2-jet invariant with respect to a nontrivial action of $\mathbb{Z}_{p}$ is at least p1p-1 for prime $p$. In this paper we prove a generalization of this inequality for actions of arbitrary abelian groups using M. Roberts’ results on invariant morsifications for invariants of real group actions.
1. Introduction
In [1], M. E. Kazarian has posed a following conjecture [^1] This is essentially a local version of the Matsumura-Monsky theorem([5]). Actually the proof by Matsumura and Monsky can be easily adapted to this statement. However, Kazarian and others did not know that, and this is a good thing, as otherwise we would not have this interesting problem on Milnor numbers to wrestle with.:
Theorem 1.1.
Consider a germ of function $f:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0)$ with an isolated singularity at the origin and a compact Lie group $G$ of analytic diffeomorphisms $(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{n},0)$ such that $f(g(x))=f(x)\;\forall\;g\in G\;\forall\;x\in\mathbb{C}^{n}$. If the Taylor series of $f$ contains no monomials of degree $\leq 2$ then $G$ is discrete.
V.A. Vassiliev has given an elementary proof of this statement based on the geometry of the Newton polytope of a germ with an infinite group of symmetries in [2]. Some time later, S.P. Chulkov has given a proof ([3]) based on the following inequality for the Milnor number of $f$:
Theorem 1.2.
Consider a germ of function $f:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0)$ with an isolated singularity at the origin such that the Taylor series of $f$ contains no monomials of degree $\leq 2$. If f is invariant with respect to a nontrivial action of a group of prime order $p$ then $\mu(f)\geq p-1$.
Kazarian’s conjecture follows from this easily, as an infinite compact Lie group contains cyclic subgroups of any order.
A slightly stronger inequality ($\mu(f)\geq p+1$) can be easily proven in case of a real group action using invariant morsifications as described in [4]. These two inequalities are essentially the only known bounds on the Milnor number in terms of properties of the group action, so it is interesting to obtain a bound for the Milnor number in case of groups of composite order. Chulkov’s original proof does not generalize (as far as we know) to groups other that $\mathbb{Z}_{p}$.
The main theorem of this paper is the generalization of Chulkov’s inequality to arbitrary abelian groups. For a group action $\tau$ on $(\mathbb{C}^{n},0)$ we will denote the length of shortest orbit of a non-fixed point by $l(\tau)$.
Theorem 1.3.
Let $f$ be a germ of analytic function $(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0)$ with an isolated singularity at the origin such that the Taylor series of $f$ contains no monomials of degree $\leq 2$. Let $\tau$ be a nontrivial action of a finite abelian group $G$ on $(\mathbb{C}^{n},0)$ by analytic diffeomorphisms. If $f$ is invariant with respect to $\tau$ then $\mu(f)\geq l(\tau)-1$.
In the case $G=\mathbb{Z}_{p}$ this reduces to the original Chulkov’s inequality.
2. Definitions and auxiliary statements
In a suitable local system of coordinates any finite group action on $(\mathbb{C}^{n},0)$ can be linearized (see [6]). So we wil consider all group actions to be induced by a linear representation of $G$. Furthermore, any linear representation of a finite abelian group is diagonalazable, hence further we will consider only actions of the form $g\circ(x_{1},\dots,x_{n})\longrightarrow(\chi_{1}(g)x_{1},\dots,\chi_{n}(g)x_{n})$ with $\chi_{i}$ being group characters of $f$.
A function $f:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0)$ is called equivariant with respect to group action $\tau$ on $(\mathbb{C}^{n},0)$ if there is a character $\lambda$ of $G$ such that $f(\tau(g)\circ x)=\lambda(g)f(x)\;\forall\;x\;\forall\;g\;\in\;G$. $\lambda$ is called the weight of $f$. Invariant functions are equivariant with weight 1, of course.
For a function germ $f$ $J_{f}$ will denote the Jacobian ideal $\frac{\partial f}{\partial x_{1}},\dots,\frac{\partial f}{\partial x_{n}}>$. $Q_{f}$ is the Jacobian algebra $\mathbb{C}{x_{1},\dots,x_{n}}/J_{f}$. $Q_{f}$ always admits a monomial basis, i.e. the basis consisting of monomials
If $f$ is invariant with respect to an action of a group $G$, the Jacobian ideal is invariant and therefore $G$ also acts on $Q_{f}$ by changes of coordinates: $g\circ h(x)=h(g^{-1}(x))$. The representation of $G$ on $Q_{f}$ is called the equvariant Milnor number of $f$ and denoted $\mu_{G}(f)$. $\nu(f)$ is the multiplicity of the trivial representation in $\mu_{G}(f)$, i.e. the dimension of the space of invariant elements of $Q_{f}$. This subspace of invariant elements is denoted $Q^{G}_{f}$.
A group action on $(\mathbb{C}^{n},0)$ on is called real if it admits an invariant quadratic form of rank $n$.
The next two theorems are taken from [4].
Theorem 2.1.
If $f$ is invariant with respect to a real action of a finite group $G$ then there is a $G$-invariant deformation $f_{\lambda}$ of $f$ with only nondegenerate critical points.
The set $C$ of critical points of any invariant deformation $f_{\lambda}$ is $G$-invariant. The action of $G$ on $(\mathbb{C}^{n},0)$ induces a representation $\theta$ of $G$ on the space of $\mathbb{C}$-valued functions on $C$ ($\theta(g)\circ h(x)=h(g^{-1}(x))$.
Theorem 2.2.
If the deformation $f_{\lambda}$ has only nondegenerate critical points then the representations $\theta$ and $\mu_{G}(f)$ are isomorphic.
The last statement mean in particular that $\nu(f)$ is equal to the number of orbits of critical points in a invariant Morse approximation of $f$.
Theorem 2.3.
If the main theorem is true for group action with no fixed points outside of origin, it is true for all group actions.
Proof.
Consider the action $\tau:g\circ(x_{1},\dots,x_{n})\longrightarrow\\ \longrightarrow(x_{1},\dots,x_{k},\chi_{k+1}(g)x_{k+1},\dots,\chi_{n}(g)x_{n}),\chi_{k+1},\dots,\chi_{n}\neq 1$ and an invariant function $f$ with a zero 2jet2-jet and an isolated critical point at the origin.
[…]
Since $\nu(f)$ is equal to the number of orbits of critical points of an invariant Morse approximation, in the case of a real representation $\tau$ with no fixed points outside of origin we obtain the inequality $\mu(f)\geq(\nu(f)-1)l(\tau)+1$. We will refer to this formula as Roberts’ inequality.
3. Proof of main theorem
For a group action $\tau:g\circ(x_{1},\dots,x_{n})\longrightarrow(\chi_{1}(g)x_{1},\dots,\chi_{n}(g)x_{n})$ we will denote $\tau_{\mathbb{R}}$ the action on $\mathbb{C}^{2n}$ defined as $g\circ(x_{1},\dots,x_{n},y_{1},\dots,y_{n})\longrightarrow\\ (\chi_{1}(g)x_{1},\dots,\chi_{n}(g)x_{n},\bar{\chi}_{1}(g)y_{1},\dots,\bar{\chi}_{n}(g)y_{n})$.
$\tau_{\mathbb{R}}$ is always a real action with the invariant quadratic form $\Sigma_{i}x_{i}y_{i}$. If $\tau$ has no fixed points outside of origin then neither does $\tau_{\mathbb{R}}$.
For a function $(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C},0) ff$$\oplus f$ is the function $f(x_{1},\dots,x_{n})+f(y_{1},\dots,y_{n})$.
It is easy to see that if $f$ has an isolated singularity at the origin then so does $f\oplus f$, and $\mu(f\oplus f)=(\mu(f))^{2}$. Furthermore, if $f$ is invariant with respect to an action $\tau$ then $f\oplus f$ is invariant with respect to the action $\tau_{\mathbb{R}}$.
Theorem 3.1.
If $p_{1},\dots,p_{\mu}$ form a monomial basis for the Jacobian algebra of $f$ then $p_{i}(x)p_{j}(y),1\leq i,j\leq\mu$ form a monomial basis in the Jacobian algebra of $f\oplus f$.
Proof.
Start with an arbitrary monomial basis $q_{1}(x,y),\dots,q_{\mu^{2}}(x,y)$} for $Q_{f\oplus f}$. Each $q_{i}$ factors as as a product of monomials in $x$ and $y$: $q_{i}=q_{i1}(x)q_{i2}(y)$. For each $i$ $q_{i1}\notin J_{f}$, as otherwise $q_{i}\in J_{f\oplus f}$. Hence we have a decomposition $q_{i1}=\Sigma_{j}c_{ij}p_{j}(x)$. Similarly, each $q_{i2}$ is a linear combination of $p_{j}(y)$. Therefore $p_{i}(x)p_{j}(y)$ form a system of generators for $Q_{f\oplus f}$. As there are exactly $\mu^{2}$ of them, they form a basis.
Theorem 3.2.
$\nu(f\oplus f)\geq\mu(f)$.
Proof.
Consider a monomial basis $p_{i}(x)p_{j}(y),1\leq i,j\leq\mu$ of $Q_{f\oplus f}$ as described in the previous theorem. Each $p_{i}(x)$ is an equivariant function with respect to the action $\tau$ with some weight $\lambda$. As $G$ acts by conjugate characters on $y$, $p_{i}(y)$ is equivariant with the inverse weight $\bar{\lambda}$. Hence  $\forall\;1\leq i\leq\mu\;p_{i}(x)p_{i}(y)\in Q^{G}_{f\oplus f}$.
Now we can conclude the proof of the main theorem. As $\tau_{\mathbb{R}}$ is a real representation, we can use Roberts’ inequality for the Milnor number of $f\oplus f$:
$\mu(f\oplus f)\geq(\nu(f\oplus f)-1)l(\tau_{\mathbb{R}})+1$.
As we now know $\mu(f\oplus f)=\mu^{2}(f)$, $\nu(f\oplus f)\geq\mu(f)$ and, obviously, $l(\tau_{\mathbb{R}})=l(\tau)$. Substituting all of these into Roberts’ inequality gives us the inequality
$\mu^{2}(f)\geq(\mu(f)-1)l(\tau)+1$.
Solving this quadratic inequality for $\mu(f)$, we obtain $\mu(f)=1$(which is impossible given that the 2-jet of $f$ is zero) or $\mu(f)\geq l(\tau)-1$.
4. Tightness of the bound
The next step is to produce examples showing the tightness of this bound. The trivial example is the 1-dimensional representation: if a character $\chi$ of group $G$ has order $k$ and $G$ acts on $\mathbb{C}$ via this character, the invariant function $x^{k}$ has the lowest possible Milnor number k1k-1. However, this is unsatisfactory, as we would ideally like to prove that a)the bound is tight in arbitrarily large dimensions; b)the Milnor number is bounded by the length of the shortest orbit when there are orbits of different lengths. To show this, we present the following example.
Let $n$ be an odd positive integer, $g$ a generator of a group $\mathbb{Z}_{2^{n}+1}$, $h$ a generator of $\mathbb{Z}_{3}$, $\epsilon,\sigma$ – primitive roots of unity of order $2^{n}+1,3$, respectively .
$\mathbb{Z}_{2^{n}+1}\times\mathbb{Z}_{3}$ acts on $\mathbb{C}^{n}$ in the following manner: $g\circ(x_{1},\dots,x_{n})$
$(\epsilon^{(-2)}x_{1},\epsilon^{(4)}x_{2},\dots,\epsilon^{(-2)^{n-1}}x_{n-1},\epsilon x_{n})$, $h\circ(x_{1},\dots,x_{n})=\sigma(x_{1},\dots,x_{n})$.
The polynomial $x_{1}^{2}x_{2}+x_{2}^{2}x_{3}+\dots+x_{n-1}^{2}x_{n}+x_{n}^{2}x_{1}$ is invariant and has Milnor number $2^{n}$, the length of the shortest orbit (e.g. of the point $(0,0,\dots,0,1)$) is $2^{n}+1$, however the longest orbit (e.g. of $(1,1,\dots,1)$) has length $3(2^{n}+1)$